# Connecting Star Formation in the Milky Way and Nearby Galaxies -II. An Observationally Driven Analytical Model for Predicting Cloud-Scale Star Formation Rates

### v2

@@ Abstract
– ISM: clouds – galaxies: clusters – galaxies: star formation
@@ 2.1 Correlations from the Milky Way
Figure 1: Correlations between the total clump mass in a cloud ($M_{\mathrm{clump,tot,obs}}$), the cloud mass ($M_{\mathrm{cloud}}$) and the SFR of the cloud ($\mathrm{SFR}_{\mathrm{cloud,obs}}$) in the inner Milky Way. The red dashed line represents a linear fit to the scatter points, with $k$ and $r$ denoting the slope and the correlation coefficient. The contours show the density distribution of the scatter points, derived from a Kernel Density Estimate (KDE).
@@
$\frac{\Sigma_{\rm SFR}}{M_{\odot}~yr^{-1}~kpc^{-2}}=3.8\times 10^{-3}\left(\frac{I_{\rm 24\mu m}}{MJy~sr^{-1}}\right),$ (1)
@@ 3.2 Transition mass
Figure 2: Compare the observed SFR with the model-predicted SFR. Green points represent the observation, the black dashed line shows the linear fit to the green points, with $k$ and $r$ denoting the slope and the correlation coefficient. (a) The orange and blue points are the predictions under the constant SFH based on equation 10 and equation 3, respectively, and the shaded regions indicate the uncertainty ranges; (b) Solid and dashed lines indicate the linear fits of the predictions for the 13 different SFHs, with the red line highlighting the result for the constant SFH. The black dashed line represents the observation; (c) Same as panel (a). The red dashed line shows a fit to the predictions based on equation 14, while the cyan shaded region indicates the uncertainty in the predicted results.
@@
where $a$ and $b$ specify the transition mass range. $t$ is limited to the range [0,1]. The final relation is then given by
$M_{\mathrm{clump,tot,m}}=(1-t)\,M_{\mathrm{clump,tot,obs}}+t\,M_{\mathrm{clump,tot}},$ (14)
@@
$\mathcal{L}=|R_{\text{median}}-1|+\frac{\sigma_{R}}{R_{\text{median}}}$ (15)
@@ 3.3 Uncertainty
where $i$ denotes the $i$-th parameter in the analysis. $j$ indexes the $j$-th random sample drawn for parameter $i$ at cloud mass $M_{k}$. $\text{SFR}_{i}^{(j)}(M_{k})$ is the predicted SFR obtained when parameter $i$ is set to its $j$-th randomly sampled value, while all other parameters are held fixed at their reference values. $N$ represents the Monte Carlo sample size. Then the statistics are integrated across the cloud mass range:
@@
Parameter-specific contributions to the total uncertainty are then ranked by $\overline{\text{CV}}_{i}$ in descending order. $\overline{\text{CV}}_{i}$ is prioritized over $\overline{\Delta}_{i}$ because it directly quantifies the prediction uncertainty range—how much SFR predictions scatter due to parameter variations—whereas $\overline{\Delta}_{i}$ only measures systematic bias from a reference. In uncertainty analysis, understanding the full spread of possible outcomes (captured by $\overline{\text{CV}}_{i}$) is more critical than knowing the average offset, as symmetric distributions can show zero $\overline{\Delta}_{i}$ while still having substantial uncertainty. The scale-invariance of $\overline{\text{CV}}_{i}$ also allows consistent comparison across parameters despite the five-order magnitude range in $M_{\rm cloud}$, making it the appropriate metric for ranking parameter contributions to overall model uncertainty.
@@
where $N_{\text{p}}$ is the number of parameters. As reflected in Table.2, equation 11 contributes $\approx$31 % of the total uncertainty, due to the large scatter in the molecular cloud SFE as a function of cloud mass. In Zhou et al. (2025a), the Pearson correlation coefficient of this equation is only $\approx$0.27. The significant uncertainty introduced by equation 11 is also clearly reflected in Fig.2(a), where the orange-shaded uncertainty band is substantially wider than the blue-shaded band.
@@ 3.4 Current Limitations and Caveats
In Urquhart et al. (2022), 5007 ATLASGAL clumps are classified into four evolutionary stages (quiescent, protostellar, young stellar objects and HII regions) and find similar numbers of clumps in each stage. The ATLASGAL clumps with HII regions (HII-clumps) are considered to be in the final stage of star formation and contain fully formed embedded star clusters (Wells et al., 2022; Zhou et al., 2024d, c). Assuming that HII-clumps account for 1/4 of the total number of clumps, and that the typical star formation efficiency of clumps is $\approx$0.3 (Zhou et al., 2024c, and references therein), the fraction of stellar mass contained in the clumps is 1/4*0.3 (7.5%), and can therefore be neglected.

### v3

@@ Abstract
– galaxies: ISM – galaxies: star formation – ISM: clouds – ISM: structure – stars: formation – galaxies: star clusters: general
@@ 2.1 Correlations from the Milky Way
Figure 1: Correlations between the total clump mass in a cloud ($M_{\mathrm{clump,tot,obs}}$), the cloud mass ($M_{\mathrm{cloud}}$) and the SFR of the cloud ($\mathrm{SFR}_{\mathrm{cloud,obs}}$) in the inner Milky Way. The red dashed line represents a linear fit to the scatter points, with $k$ and $r$ denoting the slope and the correlation coefficient. The contours show the density distribution of the scatter points, derived from a Kernel Density Estimate.
@@
$\frac{\Sigma_{\rm SFR}}{\rm M_{\odot}~yr^{-1}~kpc^{-2}}=3.8\times 10^{-3}\left(\frac{I_{\rm 24\mu m}}{\rm MJy~sr^{-1}}\right),$ (1)
@@ 3.2 Transition mass
Figure 2: Compare the observed SFR with the model-predicted SFR. Green points represent the observation. The black dashed line shows the linear fit to the green points, with $k$ and $r$ denoting the slope and the correlation coefficient, respectively. (a) The orange and blue points are the predictions under the constant SFH based on equation 10 and equation 3, respectively, and the shaded regions indicate the uncertainty ranges; (b) Solid and dashed lines indicate the linear fits of the predictions for the 13 different SFHs, with the red line highlighting the result for the constant SFH. The black dashed line represents the observation; (c) Same as panel (a). The red dashed line shows a fit to the predictions based on equation 14, while the cyan shaded region indicates the uncertainty in the predicted results.
@@
where $a$ and $b$ specify the transition mass range. The parameter $t$ is limited to the range [0,1]. The final relation is then given by
$M_{\mathrm{clump,tot,m}}=(1-t)\,M_{\mathrm{clump,tot,obs}}+t\,M_{\mathrm{clump,tot}}~,$ (14)
@@
$\mathcal{L}=|R_{\text{median}}-1|+\frac{\sigma_{R}}{R_{\text{median}}},$ (15)
@@ 3.3 Uncertainty
where $i$ denotes the $i$-th parameter in the analysis, and $j$ indexes the $j$-th random sample drawn for parameter $i$ at cloud mass $M_{k}$. The quantity $\mathrm{SFR}_{i}^{(j)}(M_{k})$ denotes the predicted star formation rate obtained when parameter $i$ is set to its $j$-th randomly sampled value, while all other parameters are held fixed at their reference values. The symbol $N$ represents the Monte Carlo sample size. Then the statistics are integrated across the cloud mass range:
@@
Parameter-specific contributions to the total uncertainty are then ranked by $\overline{\text{CV}}_{i}$ in descending order. The statistical measure $\overline{\text{CV}}_{i}$ is prioritized over $\overline{\Delta}_{i}$ because it directly quantifies the prediction uncertainty range—how much SFR predictions scatter due to parameter variations—whereas $\overline{\Delta}_{i}$ only measures systematic bias from a reference. In uncertainty analysis, understanding the full spread of possible outcomes (captured by $\overline{\text{CV}}_{i}$) is more critical than knowing the average offset, as symmetric distributions can show zero $\overline{\Delta}_{i}$ while still having substantial uncertainty. The scale-invariance of $\overline{\text{CV}}_{i}$ also allows consistent comparison across parameters despite the five-order magnitude range in $M_{\rm cloud}$, making it the appropriate metric for ranking parameter contributions to overall model uncertainty.
@@
where $N_{\text{p}}$ is the number of parameters. As reflected in Table.2, equation 11 contributes $\approx$31 % of the total uncertainty, due to the large scatter in the molecular cloud SFE as a function of cloud mass (Zhou et al., 2025a). The significant uncertainty introduced by equation 11 is also clearly reflected in Fig.2(a), where the orange-shaded uncertainty band is substantially wider than the blue-shaded band.
@@ 3.4 Current Limitations and Caveats
In Urquhart et al. (2022), 5007 ATLASGAL clumps are classified into four evolutionary stages (quiescent, protostellar, young stellar objects and HII regions) and find similar numbers of clumps in each stage. The ATLASGAL clumps with HII regions (HII-clumps) are considered to be in the final stage of star formation and contain fully formed embedded star clusters (Wells et al., 2022; Zhou et al., 2024d, c). Assuming that HII-clumps account for 1/4 of the total number of clumps, and that the typical star formation efficiency of clumps is $\approx$0.3 (Zhou et al., 2024c, and references therein), the fraction of stellar mass contained in the clumps is 1/4 $\times$ 0.3 (7.5%), and can therefore be neglected.
@@ 4 Conclusion
Acknowledgements
Thanks to the referee for the detailed and constructive comments, which have significantly contributed to improving this work. Thanks to Zhiqiang Yan for the helpful discussion.